OpenStudy (hijacktrolo):

Two equations are given below: a − 3b = 16 a = b − 2 What is the solution to the set of equations in the form (a, b)? (−2, −6) (−7, −9) (−11, −9) (−12, −10) plz help im really stuck

OpenStudy (anonymous):

download photo math on your phone and it will tell you numbered math problems

OpenStudy (anonymous):

and the anwers to math problem in 20 seconds

OpenStudy (hijacktrolo):

uhhhhhhh ok

jhonyy9 (jhonyy9):

a-3b=16 a=b-2 so because you know that a=b-2 the best easy way to solve this system is the substitution method - so just substitute the a=b-2 in first equation in place of a and so you can calculi easy the b ,after this just substitute this value of b in place of b inside second equation and will get the value of a hope this is understandably sure right